# A Gentle Wakeup Call: Symmetry Breaking with Less Collision Cost

### v1

@@ 4.1 Upper Bound
4.1.1 Case 1: $n\sqrt{\mathcal{C}}\leq w_{0}$.
In this case, we show that w.h.p. AH achieves a success during the halving phase. We refer to a window with size in the range $n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ as good. We first bound the latency required to reach a good window, and then analyze the expected collision cost.
@@
Assume $n\sqrt{\mathcal{C}}\leq w_{0}$. W.h.p. in $\eta$, AH terminates at or before a good window is reached.
Lemma 2
Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. W.h.p. in $\eta$, the latency of AH is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
@@
Under AH, each packet starts with an initial window size of $w_{0}=2^{\mathcal{C}^{\epsilon}}$. By assumption, the good window size is at most $w_{0}$, and so a good window will be reached during the halving phase. The halving phase executes, and by Lemma 1, w.h.p. in $\eta$ we are guaranteed to succeed at or before a good window. The number of windows that AH executes over before reaching a good window is at most $\lg(w_{0})=\mathcal{C}^{\epsilon}$, and the sample size for each such window is $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}}))\leq d\sqrt{\mathcal{C}}C^{\epsilon}$. Therefore, the latency is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
@@
Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. W.h.p. in $\eta$, the expected collision cost for AH is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
@@
In the halving phase, by Lemma 1, w.h.p. in $\eta$ a success occurs at or before a good window, which we are guaranteed by assumption to reach since $n\sqrt{\mathcal{C}}\leq w_{0}$. Thus, we consider any window of size $w\geq n\sqrt{\mathcal{C}}$ (and no more than $w_{0}$). We upper bound the probability of a collision using Lemma 3, with $m=n$ and $p=1/w$ (since $w\geq n\sqrt{\mathcal{C}}$ and $\mathcal{C}\geq 4$, we satisfy the requirement that $p<1/m$).
Consider all the slots from the samples, and index these slots from $i\geq 1$. Let $X_{i}$ be an independent random variable where $X_{i}=1$ when the $i$-th slot contains a collision; $X_{i}=0$ otherwise. Then, by Lemma 3, we have:
$Pr(X_{i}=1)$ $\leq\frac{2(n/w)^{2}}{\left(1-(n/w)\right)}.$
Let $w^{*}\in[n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ be the size of the first good window that is reached during halving phase. To get an upper bound on the expected collision cost, we first sum up the probabilities of a collision in any of the windows from an initial size $w_{0}=2^{\mathcal{C}^{\epsilon}}$ down to $w^{*}\geq n\sqrt{\mathcal{C}}$, which requires at most $C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$ halvings. Then, we multiply by the total number of slots in all the samples until a good window is reached, and then multiply by the cost of a collision.
To start, the sum of the probabilities is as follows:
$\sum_{j=0}^{C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2}\hskip-8.0pt\frac{2\left(\frac{n}{w_{0}/2^{j}}\right)^{2}}{\left(1-\left(\frac{n}{w_{0}/2^{j}}\right)\right)}$ $=\sum_{j=0}^{C^{\epsilon}-\lg{n}-\frac{(\lg\mathcal{C})}{2}}\hskip 0.0pt\frac{2n^{2}\left(\frac{2^{j}}{2^{\mathcal{C}^{\epsilon}}}\right)^{2}}{\left(1-\frac{n2^{j}}{2^{\mathcal{C}^{\epsilon}}}\right)}$
@@
$\leq\frac{2n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{j=0}^{C^{\epsilon}-\lg{n}-\frac{(\ln\mathcal{C})}{2}}\hskip 0.0pt\frac{2^{2j}}{\left(1-\frac{1}{\sqrt{{\mathcal{C}}}}\right)}$
$\leq\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{j=0}^{C^{\epsilon}-\lg{n}-\frac{(\ln\mathcal{C})}{2}}\hskip 0.0pt2^{2j}$
$=\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\frac{4^{\mathcal{C}^{\epsilon}+1}}{(4-1)4^{\lg{n}}4^{\frac{\lg{\mathcal{C}}}{2}}}$
$<\frac{8}{\mathcal{C}}.$
We note that the second line by assumption of $n\sqrt{\mathcal{C}}\leq 2^{\mathcal{C}^{\epsilon}}$. The third line holds for any $\mathcal{C}\geq 4$.
Recall that the sample size is $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})=O(\mathcal{C}^{(1/2)+\epsilon})$. Over at most $\lg(w_{0})=O(C^{\epsilon})$ samples, the total number of slots in all the samples is $O(\mathcal{C}^{(1/2)+2\epsilon})$. Thus, letting $X=\sum_{i}X_{i}$ and using our upper bound on the probability above, the expected number of collisions is :
$E\left[X\right]$ $\leq E\left[\sum_{i=1}^{O(\mathcal{C}^{(1/2)+2\epsilon})}\hskip-8.0ptX_{i}\right]=O\left(\frac{\mathcal{C}^{(1/2)+2\epsilon}}{\mathcal{C}}\right)$
Thus, multiplying this by $\mathcal{C}$, we obtain that the expected collision cost is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
4.1.2 Case 2: $n\sqrt{\mathcal{C}}>w_{0}$.
@@
In this case, we show that w.h.p. in $\eta$ AH succeeds during the doubling phase. To begin, we show that $\mathcal{C}$ is small in $n$.
@@
Next, we show argue that a success is likely once we reach a window of size $\Theta(n)$.
@@
Assume $n\sqrt{\mathcal{C}}>w_{0}$. W.h.p. in $\eta$, in the doubling phase, AH achieves a success at or before a window of size $n$ is reached.
@@
Assume $n\sqrt{\mathcal{C}}>w_{0}$. W.h.p. in $\eta$, the latency required by AH is $O(\log^{4+(1/(2\epsilon))}n)$.
@@
Under AH, all packets start with an initial window size of $w_{0}$, and they iteratively halve the window size until a window size of $2$ (in the worst case) is reached, or until a success, whichever occurs first. Pessimistically, assume no success occurs during the halving phase. By Lemma 2, the latency incurred for the halving phase is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
After that, the doubling phase begins. W.h.p. in $\eta$, by Lemma 6, a success at or before a window of size $n$; this requires $O(\log n)$ doublings and, for each window size, we have a sample of $O(\log(w_{\mbox{\tiny cur}}))=O(\log n)$ Therefore, the latency for the doubling phase is $O(\log^{2}n)$. It follows that the sum of the latency from the halving phase and the doubling phase is:
$O(\mathcal{C}^{(1/2)+2\epsilon}\log^{2}n)$ $=O(\log^{4+(1/(2\epsilon))}n)$
since $\mathcal{C}=O(\log^{1/\epsilon}n)$ by Lemma 5.
Lemma 8
Assume $n\sqrt{\mathcal{C}}>w_{0}$. W.h.p. in $\eta$, the collision cost is $O(\log^{4+(3/(2\epsilon))}n)$.
Proof
By assumption $n\sqrt{\mathcal{C}}>w_{0}$, so we can invoke Lemma 7, which states that w.h.p. in $\eta$ the latency is $O(\log^{4+(1/(2\epsilon))}n)$. Pessimistically, every slot may contain a collision, and by Lemma 5, $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, an upper bound on the total collision cost is:
$O(\mathcal{C}\,\log^{4+(1/(2\epsilon))}n))$ $=O(\log^{4+(3/(2\epsilon))}n)$
which yields the claim.
4.1.3 Total Cost of AH
We can now prove our upper bounds on the latency and expected collision cost for AH in the static setting.
Theorem 1.1. W.h.p. AH solves the static wakeup problem with the following costs:
  1. 1.
    If  $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the latency and the expected collision cost are each $O\left(\mathcal{C}^{\frac{1}{2}+2\epsilon}\right)$.
  2. 2.
    Else, the latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ latency and the expected collision cost is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
Proof
There are two cases to consider. In the first case, $n\sqrt{\mathcal{C}}\leq w_{0}$. By Lemmas 2 and 4, w.h.p. the latency and expected collision cost for AH are each $O(\mathcal{C}^{(1/2)+2\epsilon})$, yielding (1).
In the second case, $n\sqrt{\mathcal{C}}>w_{0}$. By Lemmas, 7 and 8, w.h.p. when $n\sqrt{\mathcal{C}}>w_{0}$, AH has $O(\log^{4+(1/(2\epsilon))}n)$ latency and the collision cost is $O(\log^{4+(3/(2\epsilon))}n)$, yielding (2).
Finally, we note that the relationship between $n$ and $\mathcal{C}$— a relationship established by the adversary (recall Section 1.1)—dictates which upper bound applies. Since (1) dominates (2) when $C=\Omega(\log^{\Theta(1/\epsilon)}n)$, the result follows.
4.2 Lower Bound

### v2

@@ 4 Upper Bound
The next lemma argues that success is likely in a good window.
@@
W.h.p. in $\eta$, a success occurs if AH executes in a good window.
@@
Let $w^{*}\in[n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ (recall Section 3). The probability of a success in any fixed slot of this good window is at least:
[…]
where $d^{\prime}$ is a constant that can be made arbitrarily large depending on $d$.
We then employ Theorem 4.1, with $\delta=1/2$, to obtain:
$\leq e^{-\frac{{(1/2)}^{2}d\ln(w_{\mbox{\tiny cur}})}{(5/2)3e}}$
$=e^{-\frac{d\ln(n\sqrt{\mathcal{C}})}{30e}}$
$\leq(n\sqrt{\mathcal{C}})^{-d^{\prime}}$
$\leq\eta^{-d^{\prime\prime}}$
4.1 Case 1: $n\sqrt{\mathcal{C}}\leq w_{0}$.
In this case, the good window size is less than or equal to $w_{0}$, so we argue that AH is likely to obtain a success during the halving phase. We make use of the following known result that upper bounds the probability of a collision.
@@
Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. The expected total cost of collisions for AH is $O(\mathcal{C}^{(1/2)+\epsilon})$.
@@
[…]
$G+\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$ (1)
Upper Bounding $B_{j}$. Clearly, $B_{j}$ is upper-bounded by the cost of having a collision in every single slot of iterations $1$ through $j$. In any iteration $k$, there is a sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots for each of the at most $\ln(w_{0})$ windows; thus, over iterations $1$ through $j$, the halving phases contributes $O(j\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0}))$ slots. Additionally, in iteration $k$, the corresponding doubling phase executes $2^{k}$ samples, each consisting of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots; thus, by the sum of a geometric series, over iterations $1$ through $j$, the doubling phases contributing $O(\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})2^{j})$ slots. By the above:
$B_{j}$ $=O\left(j\,d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0})+\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})2^{j}\right)\cdot\mathcal{C}$
$=O\left(j\,\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\log(w_{0})2^{j}\right)\cdot\mathcal{C}$
$=O(C^{(3/2)+2\epsilon}\,j^{2}\,2^{j})$
since $\log(w_{0})=\log(2^{\mathcal{C}^{\epsilon}})=O(\mathcal{C}^{\epsilon})$, and $\ln(w_{\mbox{\tiny cur}})\leq\ln(2^{j}w_{0})=O(j\mathcal{C}^{\epsilon})$.
By Lemma 1, the $q=1/\eta^{c}$, where recall that $\eta=\max\{n,\mathcal{C}\}$ where we can make the constant $c>0$ as large as we wish. Thus,
$\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$ $\leq\sum_{j=1}^{\infty}(1/\eta^{c})^{j}\cdot O(C^{(3/2)+2\epsilon}\,j^{2}\,2^{j})$
$=O(1)$
where the last line follows for sufficiently large $c$.
Upper Bounding $G$. To upper bound $G$, for each slot $i$ in the first halving phase of the first iteration, let $X_{i}$ be an independent random variable where $X_{i}=1$ if the $i$-th slot contains a collision; otherwise, $X_{i}=0$. Recall that the windows of this first halving phase are repeatedly halved. For a slot in a window of size $w=w_{0}/2^{k}$, for a positive integer $k\geq 0$, by Lemma 2, with $m=n$ and $p=1/w$ (since $w\geq n\sqrt{\mathcal{C}}$ and $\mathcal{C}\geq 4$, we satisfy the requirement that $p<1/m$), we have:
$Pr(X_{i}=1)$ $\leq\frac{2(n/w)^{2}}{\left(1-(n/w)\right)}$
$=\frac{2(n/(w_{0}/2^{k}))^{2}}{\left(1-(n/(w_{0}/2^{k}))\right)}$
The expected cost due to collisions that occur in this particular window is thus at most:
@@
$Pr(X_{i}=1)\cdot d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\cdot\mathcal{C}$ $\leq\frac{2(n/(w_{0}/2^{k}))^{2}}{\left(1-(n/(w_{0}/2^{k}))\right)}d\mathcal{C}^{(3/2)+\epsilon}.$
Note that the number of windows executed starting from an initial size $w_{0}=2^{\mathcal{C}^{\epsilon}}$ down to size $n\sqrt{\mathcal{C}}$, is at most $C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$; for ease of presentation, let $x=C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$. By the above, we can derive:
$G$ $\leq\sum_{k=0}^{x}\hskip 0.0pt\frac{2\left(\frac{n}{w_{0}/2^{k}}\right)^{2}}{\left(1-\left(\frac{n}{w_{0}/2^{k}}\right)\right)}\left(d\mathcal{C}^{(3/2)+\epsilon}\right)$
$=\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\sum_{k=0}^{x}\hskip 0.0pt\frac{2n^{2}\left(\frac{2^{k}}{2^{\mathcal{C}^{\epsilon}}}\right)^{2}}{\left(1-\frac{n2^{k}}{2^{\mathcal{C}^{\epsilon}}}\right)}$
$\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{2n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt\frac{2^{2k}}{\left(1-\frac{1}{\sqrt{{\mathcal{C}}}}\right)}$ (2)
$\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt2^{2k}$ (3)
$=\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\frac{4^{\mathcal{C}^{\epsilon}+1}}{(4-1)4^{\lg{n}}4^{\frac{\lg{\mathcal{C}}}{2}}}$ (4)
$<\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{8}{\mathcal{C}}$
$=O\left(\mathcal{C}^{(1/2)+\epsilon}\right).$
In the above, line 2 holds since $n\sqrt{\mathcal{C}}\leq 2^{\mathcal{C}^{\epsilon}}$, since we are in Case 1. Line 3 holds for any $\mathcal{C}\geq 4$, and line 4 holds by the sum of the geometric series. Thus, Equation 1 is bounded as claimed.
The next lemma bounds the expected latency of our algorithm, and its argument is similar to that of Lemma 3, although an extra $O(\mathcal{C}^{\epsilon})$-factor appears. Informally, this occurs since, in counting slots for the first halving phase, we have a sample size of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})=d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $\leq d\sqrt{\mathcal{C}}\ln(w_{0})$ $d\sqrt{\mathcal{C}}C^{\epsilon}=O(\mathcal{C}^{(1/2)+\epsilon})$, and we have $\log_{2}w_{0}=O(C^{\epsilon})$ samples.
Lemma 4
Assume that $n\sqrt{\mathcal{C}}\leq w_{0}$. The expected latency of AH is $O(\mathcal{C}^{(1/2)+2\epsilon})$.
4.2 Case 2: $n\sqrt{\mathcal{C}}>w_{0}$.
@@
In this case, the good window size exceeds $w_{0}$, and so we argue that AH likely succeeds during the doubling phase. A key insight is that, for Case 2, $\mathcal{C}$ is small in terms of $n$.
@@
The next lemma establishes bounds on expected latency and total collision cost, and the intuition is as follows. In Case 2, collisions are “cheap” by Lemma 5, in the sense that $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, the cost of collisions incurred until the good window is reached is small; it takes only $\log_{2}(n\sqrt{\mathcal{C}})+O(1)=O(\log n)$ doublings of the window size to reach a good window in the doubling phase. Since $2^{j}$ counts the number of doublings we perform in the doubling phase (recall Section 3), we reach a good window by $j=O(\log\log n)$. After this point the probability of success is high, and so subsequent iterations contribute little to the expected total cost of collisions.
@@
Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected latency is $O\left(\log^{1/(2\epsilon)+4}n)\right)$.
Proof
The number of slots incurred for each individual halving phase is the sample size times the number of halvings, so upper bounded by:
$O\left(\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\right)\lg(w_{0})$ $=O\left(\sqrt{\mathcal{C}}\log(2^{j}w_{0})\log(w_{0})\right)$
$=O\left(\sqrt{\mathcal{C}}j\mathcal{C}^{2\epsilon}\right)$
$=O\left(j\log^{1/(2\epsilon)+2}n\right)$
where the last line follows by applying Lemma 5. Thus, over iterations $1$ to $j$, the number of slots contributed by the halving phases is $O(j^{2}\log^{1/(2\epsilon)+2}n)$
The number of slots in the doubling phase of the $j$-th iteration is:
$2^{j}d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $=O(2^{j}\sqrt{\mathcal{C}}\log(2^{j}w_{0}))$
$=O(2^{j}\sqrt{\mathcal{C}}\,j\mathcal{C}^{\epsilon})$
$=O(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n))$
slots, where the last line follow by Lemma 5. By the above, summing up the number of slots executed over iterations 1 to $j$ is $O(2^{j}\,j^{2}\,(\log^{1/(2\epsilon)+2}n))$.
Pessimistically, let us assume no success ever occurs during the halving phase. Focusing on the doubling phase, a good window is reached within at most $\log_{2}(n\sqrt{\mathcal{C}})+O(1)=\log_{2}n+O(\log\log n)$ doublings of the window size. Since $2^{j}$ counts the number of doublings we perform in the doubling phase (recall Section 3), we reach a good window by iteration $j=c^{\prime}\log\log n$, where $c^{\prime}>0$ is a sufficiently large constant. By the above, the total number of slots up until a good window is reached is:
$O\left(2^{j}\,j^{2}\,(\log^{1/(2\epsilon)+2}n)\right)$ $=O\left((\log n)\,(\log\log n)^{2}(\log^{1/(2\epsilon)+2}n)\right)$
$=O\left(\log^{1/(2\epsilon)+4}n\right)$
After this point the probability of success is high, by Lemma 1, and so subsequent iterations contribute little to the expected total cost of collisions. The expected latency is:
$O\left(\log^{1/(2\epsilon)+4}n\right)+\hskip-8.0pt\sum_{j=c^{\prime}\lg\lg n}^{\infty}\hskip-8.0ptO\left(\frac{1}{\eta}\right)^{j-c^{\prime}\lg\lg(n)+1}\hskip-5.0ptO\left(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n)\right)$
$=O\left(\log^{1/(2\epsilon)+4}n\right)+\hskip-8.0pt\sum_{j=c^{\prime}\lg\lg n}^{\infty}\hskip-8.0ptO\left(\frac{1}{n}\right)^{j-c^{\prime}\lg\lg(n)+1}\hskip-10.0ptO\left(2^{j}\,j\,(\log^{1/(2\epsilon)+1}n)\right)$
$=O\left(\log^{1/(2\epsilon)+4}n\right)$
which completes the argument.
The following lemma about the expected total cost of collisions is proved similarly to Lemma 6. The main difference is that, even if every slot contains a collision, the cost increases only by a $\mathcal{C}=O(\log^{1/\epsilon}n)$ factor.
@@
Assume $n\sqrt{\mathcal{C}}>w_{0}$. Then, the expected total cost of collisions is $O\left(\log^{3/(2\epsilon)+4}n\right)$.
4.3 Performance of AH
We can now prove our upper bounds on the latency and expected collision cost for AH.
Theorem 1.1. AH solves the wakeup problem with the following costs:
  1. 1.
    If  $\mathcal{C}=\Omega(\log^{\Theta(1/\epsilon)}n)$, then the expected latency is $O(\mathcal{C}^{\frac{1}{2}+2\epsilon})$ and the expected total cost of collisions is $O(\mathcal{C}^{\frac{1}{2}+\epsilon})$.
  2. 2.
    Else, the expected latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
@@ 4.3 Performance of AH
The relationship between $n$ and $\mathcal{C}$ is established by the adversary (recall Section 1.1), and there are there are two cases to consider. In Case 1, $n\sqrt{\mathcal{C}}\leq w_{0}$. By Lemmas 3 and 4, the expected latency and the expected total collision cost for AH are each $O(\mathcal{C}^{(1/2)+2\epsilon})$, yielding (1).
In the second case, $n\sqrt{\mathcal{C}}>w_{0}$. By Lemmas, 6 and 7, the expected latency and expected total cost of collisions are $O(\log^{4+\frac{1}{2\epsilon}}n)$ and $O(\log^{4+\frac{3}{2\epsilon})}n)$, respectively, yielding (2).
5 Lower Bound